OpenStudy (tootsi123):

The slope of the line whose equation is 5x + 3y = -2 is -5/3 -3/5 -2/3

OpenStudy (sweetburger):

get the equation into y=mx+b whatever value is m is your slope

OpenStudy (tootsi123):

I know that much but I want to know how to do the work. Like someone help me through the steps.

OpenStudy (sweetburger):

Start by putting 5x on the other side of the equation

OpenStudy (tootsi123):

Okay so I have 3y=-2-5x then I divide...by 3 giving me y=-2/3-5/3x am i correct

OpenStudy (sweetburger):

yes, excellent work :)

OpenStudy (sweetburger):

You would sometimes want to rewrite it as y= -5/3x -2/3.

OpenStudy (tootsi123):

Phew okay i didn't know if I was doing the work right or not so i wanted to see someone do it or walk me through. Thank you.

OpenStudy (mathmale):

BUT: It's very important that you enclose the right side, (2-5x), in parentheses, to indicate precisely what you are dividing. also, instead of typing the first term (above) as -5/3x, enclose that -5/3 inside parentheses.

OpenStudy (tootsi123):

Okay thank you this helps to :)